EVE 21-JULY-1999 shift summary

Alexander Ledovskoy (ledovsko@fnpxsr)
Wed, 21 Jul 1999 23:55:53 -0500

Shift Summary for EVE shift, 21-JULY-99

On Shift: Bruce, Sasha L.
Also: RickK, Patty, Marj, SET,

Hours of beam available at all intensities:
Hours of beam available at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13963 15:30 New shift crew takes over.
COSMICS run is in progress. TeV is down.
Estimate for beam is 17:00

DC3 is off (resting)

17:00 New estimate for beam is 19:00

17:40 Marj turns on DC3. It draws current as it did before.
Leaving this chamber off for several hours did not
help.

19:20 New estimate for beam is 20:00

20:20 New estimate for beam is 22:10

13970 20:25 KTEV_PED run started
20:30 run stopped.

========================================================================
Follow up Notes on Previous Shift

Note that the tapes half sticking out from the drives now are only half
full. For the next data run, APPEND to these tapes.

Done that but we messed it up later anyways (with fresh_start)
Tapes are out half full (or half empty, depends how you look at it)
New tapes are in.

========================================================================
Items to Follow up on Next Shift

- " On every start of the run you shell get ktevana errors and
you will do one (or even more than one) KFREND."
Run Manager.